OpenStudy (anonymous):

find the exact value of sin(3pi/8) a) -√(2-√2)/2 b) √(2-√2)/2 c) - √(2+√2)/2 d) √(2+√2)/2

OpenStudy (anonymous):

Hey to do this you would use the half angle formula

OpenStudy (anonymous):

So we know sin (3pi/4) right? use the formula \[\sin (\theta \div 2)=\sqrt{\frac{ 1-\cos2\theta }{ 2 }}\]

OpenStudy (anonymous):

So theta would be 3pi/4. Plug it into the formula above and solve

OpenStudy (anonymous):

and cos (6pi/4)=0. So we are left with \[\frac{ \sqrt{2} }{ 2 }\]

OpenStudy (anonymous):

what answer choice would that be though?

OpenStudy (anonymous):

one sec i will use a calculator

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Ok it would be D

OpenStudy (anonymous):

Sorry for the slow answer a friend of mine has been stuck in an airport for the past 7 hrs

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

Hey i see a mistake..Above i gave you the wrong formula for half angle it is really 1-cos (theta) in the sqrt. So you would get...\[\frac{ \sqrt{(1-\cos \theta)/2} }{ ? }\]

OpenStudy (anonymous):

and it would become...\[\sqrt{1-\frac{ \sqrt{2} }{ 2 }}\]

OpenStudy (anonymous):

and that boils down to...\[\sqrt{\frac{ 2-\sqrt{2} }{ 2 }}\]

OpenStudy (anonymous):

Which is D.

OpenStudy (anonymous):

I mean B lol

OpenStudy (anonymous):

it was D though

OpenStudy (anonymous):

yea it's D my bad..just double checked....anyway i g2g now